### v1

# QGFN: Controllable Greediness with Action Values

Abstract
Generative Flow Networks (GFlowNets; GFNs) are a family of reward/energy-based generative methods for combinatorial objects, capable of generating diverse and high-utility samples. However, biasing GFNs towards producing high-utility samples is non-trivial. In this work, we leverage connections between GFNs and reinforcement learning (RL) and propose to combine the GFN policy with an action-value estimate, $Q$, to create greedier sampling policies which can be controlled by a mixing parameter. We show that several variants of the proposed method, QGFN, are able to improve on the number of high-reward samples generated in a variety of tasks without sacrificing diversity.
Machine Learning, ICML
1 Introduction
Generative Flow Networks (Bengio et al., 2021a, b), also known as GFlowNets, or GFNs, were recently introduced as a novel generative framework in the family of energy based models (Malkin et al., 2022b; Zhang et al., 2022). Given some energy function $f(x)$ over objects $\mathcal{X}$, the promise of GFNs is to train a sampler $p_{\theta}$ such that at convergence $p_{\theta}(x)\propto\exp(-f(x))$; $\exp(-f(x))$ is also referred to as the reward $R(x)$ in GFN literature, inheriting terminology from Reinforcement Learning (RL). GFNs achieve this sampling via a constructive approach, treating the creation of some object $x$ as a sequential additive process (rather than an iterative local process à la Markov chain Monte Carlo (MCMC)). The main advantage of a GFN is its ability to generate a greater diversity of low-energy/high-reward objects compared to approaches based on MCMC or RL (Bengio et al., 2021a; Jain et al., 2022, 2023), or even Soft-RL (which, while related to GFNs, accomplishes something different by default, as discussed in Tiapkin et al., 2023).
An important control knob in GFNs has been the use of a temperature (hyper)parameter. To get more interesting samples, it is common to train a model to sample in proportion to $R(x)^{\beta}$; $\beta$ is an inverse temperature, typically $\gg 1$, which pushes the model to become greedier. Tuning this hyperparameter is non-trivial, making it challenging to train some GFNs; for example, trajectory balance (Malkin et al., 2022a) is sensitive to the “peakiness” of the reward landscape (Madan et al., 2023). Although it is possible to train temperature-conditional models (Kim et al., 2023), doing so requires a neural network to essentially learn a whole family of GFNs–no easy task, albeit doable, e.g., in multiobjective settings (Jain et al., 2023).
In this work, we propose a simple yet effective approach that allows selecting arbitrary greediness at inference time, which preserves the generality of temperature-conditionals, but can be trained in an easier way. We do so without the cost of learning a complicated family of functions and without conditionals, instead only training two models: a GFlowNet and an action-value function $Q$ (Watkins & Dayan, 1992; Mnih et al., 2013).
Armed with the forward policy of a GFN, $P_{F}$ (which decides a probability distribution over actions given the current state), and the action-value, $Q$, we show that it is possible to create a variety of controllably greedy sampling policies, which can be controlled by parameters that require no retraining. We show that it is possible to simultaneously learn $P_{F}$ and $Q$, and in doing so, to generate more high-reward yet diverse object sets. In particular, we introduce and benchmark three specific variants of our approach, which we call QGFN: $p$-greedy, $p$-quantile, and $p$-of-max.
We evaluate the proposed methods on 5 standard tasks used in prior GFN works: the fragment-based molecular design task introduced by Bengio et al. (2021a), 2 RNA design tasks introduced by Sinai et al. (2020), a small molecule design task based on QM9 (Jain et al., 2023), as well as a bit sequence task from Malkin et al. (2022a); Shen et al. (2023). The proposed method achieves higher average rewards and discovers modes more efficiently, sometimes by a large margin, than strong baselines.
We conduct an analysis of the proposed methods, investigate key design choices, and probe the methods to understand why they work. We also investigate other possible combinations of $Q$ and $P_{F}$–again, entirely possible at inference time, by combining trained $Q$ and $P_{F}$ models.
2 Background
We follow the general setting of previous GFN literature and consider the generation of discrete finite objects, but note that in principle our method could be extended to the continuous case (Lahlou et al., 2023).
GFlowNets
For a thorough treatment of GFNs, see Bengio et al. (2021b). GFNs sample objects by decomposing their generation process in a sequence $\tau=(s_{0},..,s_{T})$ of constructive steps. The space can be described by a pointed directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{S},\mathcal{A})$, where $s\in\mathcal{S}$ is a partially constructed object, and $(s\to s^{\prime})\in\mathcal{A}\subset\mathcal{S}\times\mathcal{S}$ is a valid additive step (e.g., adding a fragment to a molecule). $\mathcal{G}$ is rooted at a unique initial state $s_{0}$.
GFNs are trained by pushing a model to satisfy so-called balance conditions of flow, whereby flows $F(s)$ going through states are conserved such that terminal states (corresponding to fully constructed objects) are sinks that absorb $R(s)$ (non-negative) units of flow, and intermediate states have as much flow coming into them (from parents) as flow coming out of them (to children). This can be described succinctly as follows, for any partial trajectory $(s_{n},..,s_{m})$:
$F(s_{n})\prod_{i=n}^{m-1}P_{F}(s_{i+1}|s_{i})=F(s_{m})\prod_{i=n}^{m-1}P_{B}(s_{i}|s_{i+1})$ (1)
where $P_{F}$ and $P_{B}$, the forward and backward policies, are distributions over children and parents respectively, representing the fraction of flow emanating forward and backward from a state. By construction for terminal (leaf) states $F(s)=R(s)$. Another way to think of the above is via edge flows (Bengio et al., 2021a), $F(s\to s^{\prime})=F(s)P_{F}(s^{\prime}|s)$. Again, flow conditions preserve incoming flows and outgoing flows for all states $s\in\mathcal{S}$:
$\sum_{s_{i}\in\mathrm{Parents}(s)}F(s_{i}\to s)=\sum_{s_{o}\in\mathrm{Children}(s)}F(s\to s_{o})$ (2)
where, by construction, the edge flow $F(s\to s_{T})$ to a terminal (sink) state $s_{T}$ is $R(s_{T})$ (e.g. the flow corresponding to taking the stop action), and the initial state $s_{0}$ (which has no parents) only has to account for the flow of its children (it is a source in the network).
Flow conditions lead to learning objectives such as Trajectory Balance (TB; Malkin et al., 2022a), where $n=0$ and $m$ is the trajectory length, and Sub-trajectory Balance (SubTB; Madan et al., 2023), where all combinations of $(n,m)$ are used (in a way reminiscent of TD($\lambda$); Sutton, 1988). While a variety of GFN objectives exist, we use these two standard baselines. If flow conditions are fully satisfied, i.e. 0-loss everywhere, terminal states are guaranteed to be sampled with probability $\propto R(s)$ (Bengio et al., 2021a).
Note that, by construction, the relationship between $F$ and $P_{F}$ is such that, if $s\to s^{\prime}\in\mathcal{A}$, then $P_{F}(s^{\prime}|s)=P_{B}(s|s^{\prime})F(s^{\prime})/F(s)\propto F(s^{\prime})$; the likelihood of going from $s$ to $s^{\prime}$ is proportional to the flow in $s^{\prime}$.
Action values
For a broad overview of RL, see Sutton & Barto (2018). A central object in RL is the action-value function $Q^{\pi}(s,a)$, which estimates the expected “reward-to-go” when following a policy $\pi$ starting in some state $s$ and taking action $a$; for some discount $0\leq\gamma\leq 1$,
$Q^{\pi}(s,a)=\mathbb{E}_{\begin{subarray}{c}a_{t}\sim\pi(.|s_{t})\\ \!\!\!s_{t+1}\sim T(s_{t},a_{t})\end{subarray}}\left[\sum^{\infty}_{t=0}\gamma^{t}R(s_{t})\big{|}s_{0}=s,a_{0}=a\right]$ (3)
Note that $T(s,a)$ is a stochastic transition operator, but that in a GFN context objects are typically constructed in a deterministic setting (although there are stochastic extensions; Pan et al., 2023). Because rewards are only available for finished objects, $R(s)=0$ unless $s$ is a terminal state, and we use $\gamma=1$ to avoid arbitrarily penalizing “larger” objects. Finally, as there are several possible choices for $\pi$, we will simply refer to $Q^{\pi}$ as $Q$ when statements apply to a large number of such choices.
3 Motivation
Consider the following scenario, illustrated in Figure 1: an agent is faced with two doors. Behind the left door, there are 100 other doors, each hiding a reward of 1. Behind the right door, there is a single door hiding a reward of 100. The flow will be such that $F(\mathrm{left})=F(\mathrm{right})=100$, meaning that a GFN agent will pick either door with probability $\nicefrac{{1}}{{2}}$. On the other hand, the action value function is $Q(s_{0},\mathrm{left})=1$, $Q(s_{0},\mathrm{right})=100$, so an agent basing its decisions on $Q$ will reach the door with reward 100.
Figure 1: Illustration showing why solely relying on flow functions $F$ in GFNs can be insufficient. While GFNs capture how much stuff there is, they spend time sampling from lots of small rewards.
The above example shows that relying solely on flows is not always sufficient to provide high-value samples frequently and is illustrative of real-world scenarios. Consider molecular design (a very large search space) with some goodness reward in $[0,1]$; there may be $10^{6}$ molecules with reward .9.9.9.9, but just a dozen with reward $1$. Since .$9\times 10^{6}.$ is much bigger than $12\times 1$, the probability of sampling a reward 1 molecule will be low if one uses this reward naively. While using a temperature parameter is a useful way to increase the probability of the reward $1$ molecules, we propose a complementary, inference-time-adjustable method.
However, relying solely on $Q$ is also insufficient. If $Q$ were estimated very well for the optimal policy (which is extremely hard), it would be (somewhat) easy to find the reward 1 molecules via some kind of tree search following $Q$ values. However, in practice, RL algorithms easily collapse to non-diverse solutions, only discovering a few high reward outcomes. This is where flows are useful: because they capture how much stuff there is in a particular branch (rather than an expectation), it is useful to follow flows to find regions where there is potential for reward. In this paper we propose a method that can be greedier (by following $Q$) while still being exploratory and diverse (by following $F$ through $P_{F}$).
A commonly used trick in GFNs to increase the average reward, aka “greediness” of the generated samples, is to adapt the reward distribution by using an altered reward function $\hat{R}(x)=R(x)^{\beta}$ and adjusting the exponent parameter $\beta$: the higher the $\beta$, the greedier the model should be (Jain et al., 2023). However, increasing $\beta$ induces greater numerical instability. Even on a log scale, using e.g. $\beta=10$ scales a mean squared loss more or less 100×100$\times$100 ×. Another common pitfall is that increasing $\beta$ reduces the diversity of the model, because the model is less incentivized to explore “middle-reward” regions. This can lead to mode collapse. Kim et al. (2023) show that it is possible to train models that are conditioned on $\beta$, which somewhat alleviates these issues, but at the cost of training a more complex model, possibly restricted to values of $\beta$ seen during training.
Another possible approach is to leverage the equivalence between the GFN framework and the Soft-RL framework (Tiapkin et al., 2023). However, this approach produces a soft policy, while we are interested in increasing the greediness of the policy.
4 QGFN: controllable greediness through $Q$
Leveraging the intuition from the example above, we now propose and investigate several ways in which we can use $Q$-functions to achieve our goal; we call this general idea QGFN. In particular, we present three variants of this idea, which are easy to implement and effective: $p$-greedy QGFNs, $p$-quantile QGFNs, and $p$-of-max QGFNs. In §5 and §6, we show that these approaches provide a favourable trade-off between reward are diversity of the generated samples, during both training and inference.
As is common in GFlowNets, we train QGFN by sampling data from some behavior policy $\mu$. We train $F$ and $P_{F}$ (and use a uniform $P_{B}$) to minimize a flow consistency loss on the minibatch of sampled data. Additionally, we train a $Q$-network to predict action values on the same minibatch (the choice of loss will be detailed later). Training the GFN and $Q$ on a variety of behaviors $\mu$ is possible because both are off-policy methods. Indeed, instead of choosing $\mu$ to be a noisy version of $P_{F}$ as is usual for GFNs, we combine the predictions of $P_{F}$ and $Q$ to form a greedier behavior policy. In all proposed variants, this combination is modulated by a factor $p\in[0,1]$, where $p=0$ means that $\mu$ depends only on $P_{F}$, and $p=1$ means $\mu$ is greediest, as reflected by $Q$. The variants differ in the details of this combination.
$p$-greedy QGFN
Here, we define $\mu$ as a mixture between $P_{F}$ and the $Q$-greedy policy, controlled by factor $p$:
$\mu(s^{\prime}|s)=(1-p)P_{F}(s^{\prime}|s)+p\mathbb{I}[s^{\prime}=\mathrm{argmax}_{i}Q(s,i)]$ (4)
In other words, we follow $P_{F}$, but with probability $p$, the greedy action according to $Q$ is picked. All states reachable by $P_{F}$ are still reachable by $\mu$. Note that $p$ can be changed to produce very different $\mu$ without having to retrain anything.
$p$-quantile QGFN
Here, we define $\mu$ as a masked version of $P_{F}$, where actions below the $p$-quantile of $Q$, denoted $q_{p}(Q,s)$, have probability 00 (so are discarded):
$\mu(s^{\prime}|s)\propto P_{F}(s^{\prime}|s)\mathbb{I}[Q(s,s^{\prime})\geq q_{p}(Q,s)]$ (5)
This can be implemented by sorting $Q$ and masking the logits of $P_{F}$ accordingly. This method is more aggressive, since it prunes the search space, potentially making some states unreachable. Again, $p$ is changeable.
$p$-of-max QGFN
Here, we define $\mu$ as a masked version of $P_{F}$, where actions with $Q$-values less than $p\max_{a}Q(s,a)$ have probability 0:
$\mu(s^{\prime}|s)\propto P_{F}(s^{\prime}|s)\mathbb{I}[Q(s,s^{\prime})\geq p\max_{i}Q(s,i)]$ (6)
This is similar to $p$-quantile pruning, but the number of pruned actions changes as a function of $Q$. If all actions are estimated as good enough, it may be that no action is pruned, and vice versa, only the best action may be retained is none of the others are good.. This method also prunes the search space, and $p$ remains changeable. Note that in a GFN context, rewards are strictly positive, so $Q$ is also positive.
Policy evaluation, or optimal control?
In the design of the proposed method, we were faced with an interesting choice: what policy $\pi$ should $Q^{\pi}$ evaluate? The first obvious choice is to perform $Q$-learning (Mnih et al., 2013), and estimate the optimal value function $Q^{*}$, with a 1-step TD objective. As we detail in §6, this proved to be fairly hard, and in general, the learned $Q_{\theta}$ ended up being a poor approximation.
A commonly used trick to improve the performance of bootstrapping algorithms is to use $n$-step returns (Hessel et al., 2018). This proved essential to our work, and also revealed something curious: we’ve consistently found that, while results started improving at $n\geq 5$, a consistently good value of $n$ was close to the maximum trajectory length. This has an interesting interpretation, as beyond a certain value of $n$, $Q$ becomes closer to $Q^{\mu}$ and further from $Q^{*}$. In other words, using an “on-policy” estimate $Q^{\mu}$ rather than an estimate of the optimal policy seems beneficial, or at least easier to learn as a whole. In hindsight, this makes sense because on-policy evaluation is easier than learning $Q^{*}$, and since we are combining the $Q$-values with $P_{F}$, any method which biases $\mu$ correctly towards better actions is sufficient (we do not need to know exactly the best action, or its exact value).
Selecting greediness
In the methods proposed above, $p$ can be changed arbitrarily. We first note that we train with a constant or annealed [^1] Specifically for $p$-quantile QGFN and $p$-of-max, we found training to be more stable if we started with $p=0$ and annealed $p$ towards its final value with a single $\nicefrac{{1}}{{2}}$-period cosine schedule over 1500 steps. value of $p$ and treat it as a hyperparameter. Second, as discussed in §6, after training, $p$ can be changed with a predictable effect: the closer $p$ is to 1, the greedier $\mu$ becomes. Presumably, this happens because the model is generalizing, and $Q$-value estimates for “off-policy” actions are still a reliable guess of the reward obtainable down some particular branch. When making $p$ higher, $Q$ may remain a good lower bound of the expected reward (after all, $\mu$ is becoming greedier), which is still helpful. Generally, such a policy will have reasonable outcomes, regardless of the specific $\mu$ and $p$ used during training Finally, we note it may be possible, and desirable, to use more complex schedules for $p$, or to sample $p$ during training from some (adaptive) distribution.
5 Main results
We experiment on 4 standard tasks, detailed below, which were used in prior GFlowNet literature. As baselines, we use Trajectory Balance, Sub-Trajectory Balance, LSL-GFN (Kim et al., 2023) i.e. learning to scale logits which controls greediness through temperature-conditioning, as RL baselines A2C (Mnih et al., 2016) and DQN (Mnih et al., 2013) (which on its own systematically underperforms in these tasks).
We report the average reward obtained by the agents, as well as the total number of modes of the distribution of interest found during training. By mode, we mean a high-reward object that is separated from previously found modes by some distance threshold. The distance function and threshold we use, as well as the minimum reward threshold for an object to be considered a mode, depend on the task.
Fragment-based molecule generation task: Generate a graph of up to 9 fragments, where the reward is based on a prediction of the binding affinity to the sEH protein, using a model provided by Bengio et al. (2021a). $|\mathcal{X}|\approx 10^{100}$, there are 72 available fragments, some with many possible attachment points. We use Tanimoto similarity (Bender & Glen, 2004), with a threshold of 0.65, and a reward threshold of 0.97. Results are shown in Fig. 2.
Atom Based QM9 task: Generate small molecules of up to 9 atoms following the QM9 dataset (Ramakrishnan et al., 2014). $|\mathcal{X}|\approx 10^{12}$, the action space includes adding atoms or bonds, setting node or bond properties and stop. A MXMNet proxy model (Zhang et al., 2020), trained on QM9, predicts the HOMO-LUMO gap, a key indicator of molecular properties including stability and reactivity, and is used as the reward. Rewards are in the $[0,2]$ range, with a $1.10$ threshold and a minimum Tanimoto similarity of $0.70$ to define modes. Results are shown in Fig. 3.
RNA-binding task: Generate a string of 14 nucleobases. The reward is a predicted binding affinity to the target transcription factor, provided by the ViennaRNA package for the binding landscapes (Lorenz et al., 2011); we experiment with two RNA binding tasks; L14-RNA1, and L14-RNA1+2 (two binding targets) with optima computed from Sinai et al. (2020). $|\mathcal{X}|$ is $4^{14}\approx 10^{9}$, there are 4 tokens: adenine (A), cytosine (C), guanine (G), uracil (U). Results are shown in Fig. 4.
Prepend-Append bit sequences: Generate a bit sequence of length 120 in a prepend-append MDP, where $|\mathcal{X}|$, limited to $\{0,1\}^{n}$, is $2^{120}\approx 10^{36}$. For a sequence of length $n$, $R(x)=\exp\left(1-\min_{y\in M}d(x,y)/n\right)$. A sequence is considered as a unique mode if it is within the edit distance threshold $\delta$ from $M$, where $M$ is defined as per Malkin et al. (2022a) (it is also worth noting that the task we consider here is a more complex version, introduced by Shen et al. (2023), since prepend actions induce a DAG instead of a simpler tree). In our experiment, $|M|=60,n=120,k=1,\delta=28$. Results are shown in Fig. 5.
Figure 2: Fragment-based molecule generation task. Left: Average rewards over the training trajectories. Center: Number of unique modes with a reward threshold exceeding 0.97 and pairwise Tanimoto similarity score less than 0.65. Right: Average pairwise Tanimoto similarity score for the top 1000 molecules sampled by reward. Lines are the interquartile mean and standard error calculated over 5 seeds.
Figure 3: QM9 task. Left: Average rewards over training trajectories. Right: Number of modes with a reward above 1.10 and pairwise Tanimoto similarity less than 0.70.
Figure 4: RNA-binding tasks: Left: Average reward and modes for L14RNA1 task. Right: Average reward and modes for L14RNA1+2 task, based on 5 seeds (interquartile mean and standard error shown).
Figure 5: Bit sequence generation task. Interquartile mean and standard error of with $k=1$ over 5 seeds.
5.1 Analysis of results
Across tasks, QGFN variants produce higher rewards and find a higher number of modes, i.e. high-reward dissimilar objects. This could seem surprising, because a priori, increasing the greediness of a method likely reduces its diversity. This fundamental trade-off is known in RL as the exploration-exploitation dilemma (Sutton, 1988; Sutton & Barto, 2018). However, we are leveraging two methods and combining their strengths to reap the best from both worlds: GFlowNets are able to cover the state space, because they attempt to model all of it, by learning $p_{\theta}(x)\propto R(x)$, while $Q$ approximates the expected reward of a particular action, which can help guide the agent by selecting high-expected-reward branches. Another way to think about this: GFNs are able to estimate how many high-reward objects there are in different parts of the state space. The agent thus ends up going in all important regions of the state space, but by being a bit more greedy through $Q$, it focuses on higher reward objects, so it is more likely to find objects with reward past the mode threshold.
We also report the average reward and pairwise similarity based on 1000 samples taken after training in Table 1. Again, QGFNs outperform GFNs in reward, while retaining low levels of inter-sample similarity. We also note that at inference, we are able to use a different (and better) $p$ value than the one used at training time. We expand on this in §6, and show that it is easy to tune $p$ to achieve different reward-diversity trade-offs at inference.
Table 1: Fragment-based Molecule Generation Task: Reward and Diversity at inference with pre-trained models. The $p$ values for $p$-greedy, $p$-of-max, and $p$-quantile QGFN are set at 0.4, 0.9858, and 0.93, respectively.
 Method Reward($\uparrow$) Similarity ($\downarrow$)
 GFN-TB 0.780$\pm$0.003 0.545$\pm$0.002
 GFN-SubTB 0.716$\pm$0.006 0.513$\pm$0.003
 p-greedy QGFN 0.950$\pm$0.004 0.551$\pm$0.015
 p-of-max QGFN 0.969$\pm$0.003 0.514$\pm$0.001
 p-quantile QGFN 0.955$\pm$0.003 0.509$\pm$0.008
QGFN variants matter
We point the reader to an interesting result, which is consistent with our understanding of the proposed method. In the fragment task, the number of actions available to the agent is quite large, ranging from about 100-1000 actions, and the best performing QGFN variant is one that consistently masks most actions: $p$-quantile QGFN. It is likely indeed that most actions are harmful, as combining two fragments that do not go together may be irreversibly bad, and masking helps the agent avoid undesirable regions of the state space. However, masking a fixed ratio of actions can provide more stable training.
[…]
6 Method Analysis
We now analyze the key design choices in the methods we presented. We start by investigating the impact of $n$ (the number of bootstrapping steps in $Q$-Learning) and $p$ (the mixture parameter) on training. We then look at trained models, and reuse the learned $Q$ and $P_{F}$ to show that it is possible to use a variety of sampling strategies, and to change the mixture factor $p$ to obtain a spectrum of greediness. Finally, we run analysis experiments to provide evidence as to why QGFN is helpful.
Impact of $\beta$ in QGFN
Fig. 6 shows the effect of training with different $\beta$ values on the average reward and number of modes when taking 1000 samples after training is done in the fragment task (over 5 seeds). As predicted, increasing $\beta$ increases the average reward of the agent, but at some point, causes it to become dramatically less diverse. As discussed earlier, this is typical of GFNs with a too high $\beta$, and is caused by a collapse around high-reward points and an inability for the model to further explore. While QGFN is also affected by this, it does not require as drastic values of $\beta$ to obtain a high average reward.
Figure 6: Effect of $\beta$; fragment task. Increasing greediness through $\beta$ increases the average reward, but can cause diversity collapse. QGFN maintains diversity with a lower $\beta$, while GFN collapses. Modes are counted from 1000 samples at inference, using an inference-adjusted $p$.
Figure 7: Effect of training $p$, and $n$; fragment task. Changing $p$ can control greediness. Increasing $n$ is generally beneficial. Modes are counted from 1000 samples generated using the training $p$.
Impact of $n$ in QGFN
As mentioned in §4, training $Q$ with 1-step returns is ineffective and produces less useful approximations of the action value. Fig. 7 shows the number of modes from 1000 post-training samples in the fragment tasks, for models trained with a variety of $n$-step values. Models start being consistently good at $n=5$ and values close to the maximum length of a trajectory tend to work well too.
Impact of the training $p$ in QGFN
While our method allows changing $p$ more or less at will, we still require some value during training. Fig. 7 shows that there are clear trade-offs between choices of $p$, some yielding significantly better diversity than others. For example, $p$-of-max is fairly sensitive to the chosen value during training, and for the fragment task doesn’t seem to perform particularly well during training. On the other hand, as we will see in the next paragraph (and is also seen in Fig. 6), $p$-of-max excels at inference, and is able to generate diverse and high-reward samples by adjusting $p$.
Changing strategy after training
We now look at the impact of changing the mixture parameter $p$ and the sampling strategy for already trained models on average reward and average pairwise similarity. We use the parameters of a model trained with $p$-greedy QGFN, $p=0.4$.
With this model, we sample 512 new trajectories for a series of different $p$ values. For $p$-greedy and $p$-quantile, we vary $p$ between 00 and $1$; for $p$-of-max, we vary $p$ between .9.9.9.9 and $1$ (values below .9.9.9.9 have minor effects). We visualize the effect of $p$ on reward and similarity statistics in Fig. 8.
First, we note that increasing $p$ has the effect we would hope, increasing the average reward. Second, we note that this works without any retraining; even though we (a) use values of $p$ different than those used during training, and (b) use QGFN variants different than those used during training, the behavior is consistent: $p$ controls greediness. Let us emphasize (b): even though we trained this $Q$ with $p$-greedy QGFN, we are able to use the $Q(s,a)$ predictions just fine with entirely different sampling strategies. This has some interesting implications; most importantly, it can be undesirable to train with too high values of $p$ (because it may reduce the diversity to which the model is exposed), but what is learned transfers well to sampling new, high-reward objects with different values of $p$ and sampling strategies.
Figure 8: Varying $p$ at inference time induces reward-diversity trade-offs; fragment task.
Finally, these results suggest that we should be able to prototype new QGFN variants, including expensive ones (e.g. MCTS) without having to retrain anything. We illustrate the performance of a few other variants in §E.1, Fig. 12.
Is $Q$ calibrated?
For our intuition of why QGFN works to really pan out, $Q$ has to be accurate enough to provide useful guidance towards high-reward objects. We verify that this is the case with the following experiment. We take a trained QGFN model ($p$-greedy, $p=0.4$, fragment task, maximum $n$-step) and sample $64$ trajectories. For each of those trajectories, we take a random state within the trajectory as a starting point, thereafter generating $512$ new trajectories. We then use the reward of those $512$ trajectories as an empirical estimate $\hat{Q}$ of the expected return, which $Q$ should roughly be predicting. Fig. 9 shows that this is indeed the case. Although $Q$ is not perfect, and appears to be underestimating $\hat{Q}$, it is making reasonable predictions.
Figure 9: Comparing $Q(s,a;\theta)$ predictions with empirical estimates obtained by rollouts. Bars are standard error. $Q$ is relatively capable to estimate the returns of the corresponding policy.
Is $Q$ really helpful?
In this experiment, we verify our intuition that pruning based on $Q$-values is helpful. We again take a trained model for the fragment task, and sample $512$ trajectories. We use $p$-of-max QGFN ($p=0.95$), and compare it to a strategy estimating best pruned actions: for each trajectory, after some random number of steps $t\sim U[4,20]$ (the max is 27), we start deterministically selecting the action that is the most probable according to $P_{F}$ but would be masked according to $Q$. To ensure that this is a valid thing to do, we also simply look at Best actions, i.e. after $t\sim U[4,20]$ steps, deterministically select the action that is the most probable according to $P_{F}$, regardless of $Q$.
Fig. 10 shows that our sanity check, Best actions, receives reasonable rewards, while selecting actions that would have been pruned leads to much lower rewards. The average likelihood from $P_{F}$ of these pruned actions was .035.035.035.035, while the average number of total actions was $\approx 382$ (and $1/382\approx 0.0026$). This confirms our hypothesis that $Q$ indeed masks actions that are likely according to $P_{F}$ but that do not consistently lead to high rewards.
Figure 10: Pruning helps avoid low-reward parts of the state space. Reward distributions obtained when (a) normal sampling with $p$-of-max QGFN; (b) greedily selecting actions that $p$-of-max QGFN would prune, but are the most likely according to $P_{F}$, Best pruned actions; (c) selecting most likely $P_{F}$ actions regardless of $Q$, Best actions; and (d) normal sampling from $P_{F}$ (without using $Q$).
Why does changing $p$ work?
Recall that for QGFN to be successful, we rely on $n$-step TD, and therefore on somewhat “on-policy” estimates of $Q^{\mu}$. $\mu$ is really $\mu_{p}$, a function of $p$, meaning that if we change $p$, say to $p^{\prime}$, during inference, $Q^{\mu_{p}}$ is not an accurate estimate of $Q^{\mu_{p^{\prime}}}$. If this is the case, then there must be a reason why it is still helpful to prune based on $Q^{\mu_{p}}$ while using $\mu_{p^{\prime}}$. In Fig. 11,we perform the same measurement as in Fig. 9, but we change the $p$ value used to measure $\hat{Q}^{\mu_{p^{\prime}}}$. We find that, while the rank correlation drastically goes down (although it stays well above 0), $Q^{\mu_{p}}$ remains helpful because it lower bounds $\hat{Q}^{\mu_{p^{\prime}}}$. If we prune based on $Q^{\mu_{p}}$, then we would want it to not lead us astray, and at least make us greedier as we increase $p$. This means that if an action is not pruned, then we expect samples coming from it to be at least as good as what $Q^{\mu_{p}}(s,a)$ predicts (in expectation). This is indeed the case.
Note that reducing $p$ simply leads $\mu$ to behave more like $P_{F}$, which is still a known-to-work reward-proportional sampler, and to rely less on $Q$, whose imperfections will then have less effect anyways.
7 Conclusion and Discussion
In this paper, we showed that by jointly training GFNs and $Q$-functions, we can combine their predictions to form behavior policies that are able to sample larger numbers of diverse and high-reward objects. These policies’ mixture parameter $p$ is adjustable, even after training, to modulate the greediness of the resulting policy. We implement multiple ways of combining GFNs and $Q$s, referring to the general idea as QGFN: taking a greedy with probability $p$ ($p$-greedy QGFN), restricting the agent to the top $1-p$of actions ($p$-quantile QGFN), and restricting the agent to actions whose estimated value is at least a fraction $p$ of the best possible value ($p$-of-max QGFN).
We chose to show several variants of QGFN in this paper, because they all rely on the same principle, learning $Q$, but have different properties, which lead to better or worse behavior in different tasks. For example, pruning too aggressively on a task with a small number of actions is harmful. We also hope that by showing such a diversity of combinations of $P_{F}$ and $Q$, we encourage future work that combines GFNs and RL methods in novel and creative ways.
We also analyzed why our method works. We showed that the learned action-value $Q$ is predictive and helpful in avoiding actions that have high probability under $P_{F}$ but lower expected reward. Even when $Q$ predictions are not accurate, e.g. because we sample from a different policy than the one which $Q$ models, they provide a helpful lower bound that facilitates controllable greediness.
What didn’t work     The initial stages of this project were quite different to the presented method. Instead of combining RL and GFNs into one sampling policy, we instead trained two separate agents, one GFN and one DQN. Since both are off-policy methods, the agents were able to share the training data with each other. While we were hoping that sharing “greedy” DQN data with a GFN would make it better on high-reward trajectories, we found that this was not the case; instead, the DQN agent simply slowed down the whole method–despite trying a wide variety of tricks, see §E.3.
Future work     We highlight two straightforward avenues of future work. First, there probably exist more interesting combinations of $Q$ and $P_{F}$ (and perhaps $F$), with different properties and benefits. Second, it may be interesting to further leverage the idea of pruning the action space based on $Q$, forming the basis for some sort of constrained combinatorial optimization. By using $Q$ to predict some expected property or constraint, rather than reward, we could prune some of the action space to avoid violating constraints, or to keep some other properties below some threshold (e.g. synthesizability or toxicity in molecules).
Finally, we hope that this work helps highlight the differences between RL and GFlowNet, while adding to the literature showing that these approaches complement each other well. It is likely that we are only scratching the surface of what is possible in combining these two frameworks.
Impact Statement
The research conducted in this work is not far removed from practical applications of generative models. As such, we acknowledge the importance of considering safety and alignment in applications closely related to this work such as drug discovery, material design, and industrial optimization. We believe that research on GFlowNets may lead to models that better generalize, and in the long run may be easier to align. Another important application of GFlowNets is in the causal and reasoning domains; we believe that improving in those fields may lead to easier to understand and safer models.
Acknowledgements
The bulk of this research was done at Valence Labs as part of an internship, using computational resources there. This research was also enabled in part by computational resources provided by Calcul Québec, Compute Canada and Mila. Academic authors are funded by their respective academic institution, Fonds Recherche Quebec through the FACSAcquity grant, the National Research Council of Canada and the DeepMind Fellowships Scholarship.
The authors are grateful to Yoshua Bengio, Moksh Jain, Minsu Kim, and the Valence Labs team for their feedback, discussions, and help with baselines.
Author Contributions
The majority of the experimental work, code, plotting, and scientific contributions were by EL, with support from EB. SL helped run some experiments, baselines and plots. The project was supervised by EB, and DP and LP provided additional scientific guidance. Most of the paper was written by EB. DP, LP, and EL contributed to editing the paper.
Appendix A Experiment details: Fragment-based molecule generation
In our experimental setup, we follow the exact environment specifications and implementations detailed in Malkin et al. (2022a) with the proxy model, used for evaluating molecules, provided by Bengio et al. (2021b). We set a reward threshold of 0.97 to define modes, with a pairwise Tanimoto similarity criterion of less than $0.65$. RDKit (Landrum, 2013) is used to compare pairwise Tanimoto similarity.
Training of all models was conducted using the Adam optimizer over 10,000 iterations, with a mini-batch size of 64 and a learning rate set at $1\times 10^{-4}$. The architecture of the GFlowNet models is based on a graph attention transformer (Veličković et al., 2017), with 4 layers, 128-dimensional embeddings and 2 attention heads. The reward scaling parameter $\beta$ in $R^{\beta}$ was set to 32. To follow closely the original implementation of LSL-GFN described in Kim et al. (2023), we use $\beta\sim U^{[1,64]}$, where $U$ denotes a uniform distribution. Additionally, we define a simple Multi-layer Perceptron with a hidden size of 256 as the learnable scalar-to-scalar function for the LSL-GFN. For A2C, we use a learning rate of $1\times 10^{-4}$, a training epsilon of $1\times 10^{-2}$, and an entropy regularization coefficient of $1\times 10^{-3}$. For our SubTB baseline we use SubTB(1), i.e. all trajectories are averaged with equal weight. The GFN part of QGFNs is trained with TB.
We also ran an SAC (Haarnoja et al., 2018) baseline with different $\alpha$ values of 0.5, 0.7, 0.2, along with autotuning, and a $\gamma$ value of 0.99, but we were unable to get it to discover more than 50 modes for the same amount of training iterations and mini-batch sizes.
A.1 QGFN hyperparameters:
For all variants of QGFN, we employed a grid-search approach for hyperparameter tuning, with a focus on the parameters $p$ and $n$. In the $p$-greedy QGFN variant, we selected a value of $0.4$ for $p$ from the set $\{0.2,0.4,0.6,0.8\}$, and chose an $n$ of 25 from the set $\{1,2,4,8,16,24,25,26\}$. For the $p$-of-max QGFN variant, a $p$ of 0.91 was chosen from $\{0.2,0.4,0.6,0.8,0.9,0.91,0.93,0.95\}$, with $n$ again set at 25. To ensure stability throughout the training process, we applied cosine annealing with a single-period cosine schedule over 1500 steps. An additional threshold parameter of $1\times 10^{-5}$ was applied to the condition $p\max_{i}Q(s,i)>\text{threshold}$, to prevent the initial training phase from masking actions with very small values. We also introduced a clipping of the Q-value to a minimum of 0 to prevent instability during initial training. For the $p$-quantile QGFN, a $p$ of 0.8 was selected from $\{0.6,0.7,0.8,0.9,0.95\}$, with $n=25$. For all variants of QGFN, the DQN employed had a $\tau$ of 0.95, and random action probability of $\epsilon$ was set to 0.1.
Appendix B Experiment details: QM9
In this experiment, we follow the setup described by Jain et al. (2023), but only use the HOMO-LUMO gap as a reward signal. The rewards are normalized to fall between [0, 1], although the gap proxy may range from [1,2]. As mentioned in Section 5, the modes are computed with a reward threshold of 1.10 and a pairwise Tanimoto similarity threshold of 0.70. We employ the same architecture for all models as used in the fragment-based experiments. The training models are 5,000 iterations with a mini-batch size of 64, and $\beta$ is set to 32. RDKit (Landrum, 2013) is used to compare pairwise Tanimoto similarity. We train A2C with random action probability 0.01 chosen from $\{0.1,0.01,0.001\}$ and entropy regularization coefficient 0.001 chosen from $\{0,0.1,0.01,0.001\}$.
B.1 QGFN hyperparameters:
For all variants of QGFN, we employed an exhaustive grid-search approach for hyperparameter tuning, focusing on parameters $p$ and $n$. For $p$-greedy QGFN, we selected 0.4 for $p$ from $\{0.2,0.4,0.6\}$ and 29 for $n$ from $\{11,29,30\}$. For $p$-of-max QGFN, we chose 0.6 for $p$ from $\{0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$ and 29 for $n$ from $\{28,29,30\}$. For $p$-qunatile QGFN, we selected 0.6 for $p$ from $\{0.5,0.6,0.7,0.8\}$ and 29 for $n$ from $\{11,27,28,29\}$. Similarly to the fragment task, we implemented an additional threshold of $1\times 10^{-3}$ to $p\max_{i}Q(s,i)>\text{threshold}$ and clipped Q-values to a minimum of 0 for stability during initial training. Cosine annealing over 1500 steps is used for all variants.
Appendix C Experiment details: RNA-binding task
We follow the setup of Jain et al. (2022) but using the task introduced in Sinai et al. (2020). We use a sequence transformer (Vaswani et al., 2017) architecture with 4 layers, 64-dimensional embeddings, and 2 attention heads. The training for this task is 5000 iterations over mini-batch sizes of 64. The reward scaling parameter $\beta$ is set to 8 and a learning rate of $1\times 10^{-4}$ and $1\times 10^{-2}$ for $\log Z$. In this task, $\beta\sim U^{[1,16]}$ is used for LSL-GFN. Following the approach described by Sinai et al. (2020), the modes are predefined from enumerating the entire RNA landscape for L14RNA1 and L14RNA1+2 to identify local optimal through exhaustive search. ViennaRNA (Lorenz et al., 2011) is used to provide the RNA binding landscape.
C.1 QGFN hyperparameters:
For all QGFN variants, we used grid-search for tuning hyperparameters $p$ and $n$. We used $n=13$ from the set ${12,13,14}$. For $p$-greedy QGFN, $p=0.4$ was chosen from $\{0.2,0.4,0.6,0.8\}$. For $p$-of-max QGFN, $p=0.9$ was selected from $\{0.6,0.7,0.8,0.9\}$. In $p$-quantile QGFN, we tried $p=0.25$ and $0.50$, but neither performed well due to small action spaces. A stepwise scheduler set at 500 steps is applied to $p$-of-max QGFN.
Appendix D Experiment details: Bit sequence generation
The bit sequence generation task follows the same environmental setup as Malkin et al. (2022a) with $\beta$ value as 3. We generate $|M|=60$ reference sequences by randomly combining $m=15$ symbols from an initial vocabulary $H=\{00000000,11111111,11110000,00001111,00111100\}$.
Beyond the original auto-regressive generative framework in Malkin et al. (2022a), we generate sequences in a prepend-apppend fashion motivated by Shen et al. (2023). We use a Transformer (Vaswani et al., 2023) architecture for all experiments with rotary position embeddings, 3 hidden layers with hidden dimension 64 across 8 attention heads. All methods were trained for 50,000 iterations with a minibatch size of 16. For trajectory balance, we use a learning rate of of $1\times 10^{-4}$ for the policy parameters and $1\times 10^{-3}$ for $\log Z$. For SubTB, we use the same hyperparameters as in Madan et al. (2023). For LSL-GFN, we use $\beta\sim U^{[1,6]}$, where $U$ denotes a uniform distribution.
D.1 QGFN hyperparameters:
For all variants of QGFN, we did a grid-search approach for hyperparameter tuning for $p$ and $n$. For k=1 where actions are limited to {0,1}, we set $n$ at 120, selected from $\{30,60,90,120\}$, across alla QGFN variants. For the p-greedy QGFN, we chose $p$ = 0.4; for the $p$-of-max QGFN, $p$ was set to 0.3, with cosine annealing applied at 500 steps. In the $p$-qunatile QGFN, we tested $p$ values of 0.25 and 0.50, but neither achieved good performance due to the binary nature of the action space.
Appendix E Additional experiments and plots
Figure 11: What happens to $Q(s,a)$ when changing $p$? We show here that while the rank correlation between $Q$ and the empirically estimated expected reward $\hat{Q}^{\mu_{p}}$ goes down when changing $p$, $Q$ remains a useful estimate in that it mostly lower bounds $\hat{Q}^{\mu}$. This means that, at worst, pruning based on the “wrong” $Q$ $p$ combination drops some high-reward objects, but does not introduce more lower-reward objects.
E.1 Trying other QGFN variants at inference
Since the only cost to trying to different “inference” variants of QGFN is to code them, we do so out of curiosity. We show the reward/diversity trade-off curves of these variants in Fig. 12, and include $p$-of-max as a baseline variant. As in Fig. 8 we take 512 samples for each point in the curves (except for MCTS which is more expensive). We try the following:
  • $p$-thresh, mask all actions where $Q(s,a)<p$;
  • soft-Q, not really QGFN, but as a baseline simply taking softmax$(Q/T)$ for some temperature $T$, which is varied as the greediness parameter;
  • soft-Q [0.5], as above but mixed with $P_{F}$ with a factor $p=0.5$ (i.e. $p$-greedy, but instead of being greedy, use the soft-$Q$ policy);
  • GFN-then-Q, for the first $Np$ steps, sample from $P_{F}$, then sample greedily (where $N$ is the maximum trajectory length);
  • MCTS, a Monte Carlo Tree Search where $P_{F}$ is used as the expansion prior and $\max_{a}Q(s,a)$ as the value of a state. Since this is a different sampling method, we run MCTS for a comparable amount of time to other variants, getting about 350 samples, and report the average reward and diversity.
We note that these are all tried using parameters from a pretrained $p$-greedy QGFN model. It may be possible for these variants to be much better at inference if the $Q$ used corresponded to the sampling policy.
Figure 12: Trying other possible QGFN variants.
E.2 Exploring weight sharing in QGFN
Figure 13: Weight sharing in $p$-greedy QGFN (p=0.4) with different layers
We explore the impact of weight sharing between $P_{F}$ and $Q$, as they learn from the same environments and training samples. This sharing learning approach could improve in efficiency and performance. Fig. 13 shows the impact of weight sharing on $p$-greedy QGFN, specifically focusing on sharing parameters across various layers of the graph attention transformer in the fragment-based task.
Unfortunately, naively summing the GFlowNet loss and the $Q$-Learning loss does not yield any improvements, and instead slows down learning. This may be due to several factors; most likely, interference between $Q$ and $P_{F}$, and a very different scales of the loss function may induce undesirable during training. A natural next stop would be to adjust the relative magnitude of the gradients coming from each loss, and to consider different hyperparameters (perhaps a higher model capacity is necessary), but we leave this to future work. Further exploration in this area could provide additional insights and potentially reduce training complexity.
E.3 Experiments that did not work
Several approaches were attempted prior to developing QGFN. These approaches involved sampling from both GFN and DQN independently and learning from the shared data. The key strategies explored include:
  • Diverse-Based Replay Buffer: This method stores batches of trajectories in the replay buffer and samples them based on a pairwise Tanimoto similarity threshold. It aims to diversify the experience replay during training.
  • Adaptive Reward Prioritized Replay Buffer: This strategy stores batches of trajectories in the replay buffer based on the rewards of the samples. In addition, we dynamically adjusts the sampling proportion between GFN and DQN based on the reward performance of the trajectories.
  • Weight Sharing: This involves sharing weights between GFN and DQN to potentially enhance the learning and convergence of the models.
  • Pretrained-Q for Greedier Actions: This method uses a pretrained DQN model for sampling trajectories, helping the GFN to be biased towards greedier actions in the early learning stages.
  • $n$-step returns: As per QGFN, using more than 1-step temporal differences can accelerate temporal credit assignment. This on its own is not enough to solve the tasks used in this work.
Fig 14 shows the performance of these approaches evaluated on fragment based molecule generation task.
Figure 14: Fragment-based molecule generation task; we showcase the performance of QGFN’s predecessor, which failed to beat baselines regardless of our attempts to improve it.
Figure 15: Bit sequence generation, $k=4$.
Figure 16: Pairwise Tanimoto Similarity scores assessing the impacts of $\beta$, training parameters $p$ and $n$ in the fragment task. Left: An increase in $\beta$ initially decreases sample similarities, followed by a gradual increase in similarity as the models get greedier through $\beta$. Center: Increase greediness does not always correlate with sample similarity trade-offs with QGFN, but at peak greediness, similarity scores rebound. Right: Increasing $n$ increases similarity among models.